OpenStudy (anonymous):

Find the derivative of f(x) = 5 divided by x at x = -1.

OpenStudy (anonymous):

The derivative is: \[\frac{ d }{ dx }\left( \frac{ 5 }{ x } \right)=-\frac{ 5 }{ x^2 }\] Now you want to insert \[x=-1\] and solve it.

OpenStudy (anonymous):

so its 5 ?

OpenStudy (anonymous):

@Tommynaiter can you help me with some more please? :)

OpenStudy (anonymous):

No it is -5. Because if you insert x=-1, \[x^2=(-1)^2=1\] So you get \[-5/1\]